# xOffense: An AI-driven autonomous penetration testing framework with offensive knowledge-enhanced LLMs and multi agent systems

### v1

@@ 6.3 Evaluation on Complex Real-World Exploitation Chains
6.3.1 Performance without Retrieval-Augmented Generation (No RAG)
@@
To assess the baseline capabilities of our proposed system xOffense in realistic offensive security scenarios, we conducted experiments on the same set of six real-world vulnerable machines as utilized in the VulnBot [9] evaluation: Victim1, Library2, Sar, WestWild, Symfonos2, and Funbox. This machine set—originally derived from the AI-Pentest-Benchmark—covers a diverse range of exploitation challenges, including misconfigurations, weak authentication, remote code execution, privilege escalation, and multi-step attack chains. By adopting this identical set, we ensure methodological consistency and enable a direct, fair comparison with prior work.
@@
The experiments were conducted in a fully autonomous mode, without any human intervention or Retrieval-Augmented Generation (RAG) support. Each target machine was tested in five independent runs, and the reported performance represents the best subtask completion rate per machine, following the AI-Pentest-Benchmark scoring methodology. Figure 3 presents the comparative results across multiple models, including VulnBot-Llama3.1-405B, VulnBot-DeepSeek-v3, their respective base models, and our proposed Qwen3-32B variants (base and finetuned).
@@
The results reveal several noteworthy patterns. First, Qwen3-32B-finetune consistently surpasses its base counterpart across all six machines, with particularly significant improvements on Victim1 (+0.55), Library2 (+0.30), and WestWild (+0.63). These gains highlight the effectiveness of domain-specific fine-tuning in strengthening the model’s exploitation reasoning and procedural robustness. Second, while VulnBot-DeepSeek-v3 remains highly competitive—achieving the highest score on Victim1 (0.83) and WestWild (0.71)—our fine-tuned Qwen3-32B achieves comparable or superior performance on most other machines, including leading results on Sar (0.58) and Funbox (0.54).
@@
Notably, performance disparities are strongly correlated with the complexity of exploitation chains. Targets such as Symfonos2 and Funbox, which demand multi-stage privilege escalation and exploitation of non-trivial service configurations, clearly benefit from the enhanced contextual reasoning introduced via fine-tuning. This observation underscores the critical role of model specialization in addressing the inherent unpredictability and diversity of real-world penetration testing environments. In summary, the No-RAG evaluation confirms that xOffense-Qwen3-32B-finetune can autonomously achieve competitive, and in some cases state-of-the-art, performance in realistic offensive security scenarios, even without external retrieval augmentation. This establishes a robust performance baseline for subsequent RAG-enhanced evaluations.
@@
Figure 3: Comparison of subtask completion rates across six real-world vulnerable machines in a No-RAG setting.
@@
Figure 4: Comparison of subtask completion rates across six real-world vulnerable machines with RAG setting.
6.3.2 Performance with Retrieval-Augmented Generation (RAG)
@@
When augmenting the evaluation with the Knowledge Repository module, a substantial shift in performance trends emerges across the six real-world exploitation targets (see Fig. 4). Compared to the baseline (without RAG), the Qwen3-32B-Finetune model demonstrates marked improvement, achieving perfect completion scores on Victim1 and WestWild (1.00) and notable gains on Library2 (+0.20) and Symfonos2 (+0.16). Similarly, moderate increases are observed for Sar and Funbox, reflecting the model’s enhanced capability to navigate multi-step attack chains when supported by targeted, contextually relevant prior knowledge.
@@
The gains are less pronounced for Qwen3-32B-Base, with performance remaining comparatively low on challenging targets such as Symfonos2 (0.13) and WestWild (0.25). This disparity underscores the role of fine-tuning in maximizing the benefits of retrieval augmentation—without alignment to domain-specific exploitation strategies, the retrieved information alone is insufficient to ensure consistent execution success. When compared against VulnBot baselines, Qwen3-32B-Finetune with RAG achieves competitive or superior results in four out of six targets, matching the best baseline performance on Library2 (0.80) and surpassing it on Victim1 and WestWild. This suggests that RAG integration not only mitigates the limitations of the base model but also allows the fine-tuned variant to close the gap—or in certain scenarios, outperform—human-assisted frameworks.
@@
These improvements can be attributed to three key factors: (1) the retriever’s ability to surface high-relevance exploitation procedures from a curated cybersecurity corpus, (2) the fine-tuned model’s capacity to integrate external information into coherent multi-step reasoning, and (3) the reduction of hallucination-driven dead ends, which are particularly detrimental in constrained exploitation environments. Collectively, these findings reinforce the notion that RAG is a critical enabler for scalable, high-fidelity automated penetration testing in complex real-world settings.
@@ 7.1 Internal validity
The fine-tuning of Qwen3-32B on a Chain-of-Thought–enriched penetration testing dataset introduces potential internal threats. Certain vulnerability classes and exploitation strategies are disproportionately represented, which may bias the model toward specific attack vectors while limiting its capacity to generalize to underrepresented scenarios. Moreover, the integration of prompting strategies and toolchains may embed implicit task-specific heuristics, raising the possibility that reported improvements partly reflect dataset artifacts rather than genuine reasoning ability. Such factors must be considered when interpreting performance gains on structured benchmarks.
@@ 7.2 External validity
The evaluation settings—AutoPenBench and AI-Pentest-Benchmark—approximate realistic penetration workflows yet cannot fully capture the heterogeneity of production-scale environments. Operational networks often exhibit greater variability in topology, non-standard configurations, active defenses, and deception mechanisms that remain absent from current benchmarks. In addition, adversarial tactics evolve over time, whereas benchmarks are necessarily static. Consequently, the generalizability of results to enterprise systems, heterogeneous infrastructures, or zero-day exploitation scenarios should be regarded with caution.
@@ 7.3 Construct validity
Task completion rate and exploitation success were employed as primary evaluation metrics. While suitable for quantifying functional effectiveness, these measures neglect other dimensions that are central to penetration testing practice. Attributes such as stealth, efficiency of resource utilization, time-to-compromise, and resilience against detection are critical to operational realism yet remain unaccounted for in the adopted benchmarks. Furthermore, binary success measures fail to capture partial progress or incremental compromise, potentially obscuring nuances in agent behavior across complex exploitation chains.
@@ 7.4 Reliability
7.5 Summary
In sum, although the reported findings provide strong evidence of the capabilities of xOffense, these validity concerns underscore the need for broader empirical validation. Expanding evaluations to encompass more diverse infrastructures, adversarially adaptive defenses, and richer performance metrics would strengthen the robustness, scalability, and practical applicability of autonomous penetration testing systems.
@@ 8 Conclusion and Future Work
This work presented xOffense, an independent, fully autonomous multi-agent framework for penetration testing, designed to address persistent limitations in existing systems such as context loss, limited reasoning continuity, and dependence on large proprietary models. By integrating a fine-tuned, mid-scale open-source LLM (Qwen3-32B) with a novel grey-box phase prompting mechanism and a purpose-built orchestration architecture, xOffense achieves accurate multi-stage decision-making and robust tool integration across the entire penetration testing lifecycle.
Our evaluation on AutoPenBench and AI-Pentest-Benchmark demonstrated that xOffense consistently outperforms both larger commercial LLMs (e.g., GPT-4o, LLaMA3-70B) and leading open-source baselines (e.g., PentestGPT, VulnBot-LLaMA3-405B). The framework attained an overall task completion rate of 72.72% and a sub-task completion rate of up to 79.17%, while successfully exploiting complex real-world targets. These results highlight that a domain-specialized, mid-scale model—when paired with targeted reasoning guidance—can match or exceed the capabilities of state-of-the-art large-scale systems, offering a cost-effective and reproducible solution for autonomous offensive security operations.
Future work will explore three main directions. First, we aim to optimize the command generation module, potentially through structured function calling, to further improve execution precision. Second, we plan to enhance the robustness of long-running process handling and strengthen the retrieval-augmented generation mechanism with automated updates from vulnerability intelligence sources such as ExploitDB and GitIngest. Third, we intend to extend xOffense’s capabilities to support advanced web and GUI interactions via browser automation, enabling it to tackle a broader range of penetration testing scenarios.
Acknowledgement
This research is funded by Vietnam National University Ho Chi Minh City (VNU-HCM) under grant number NCM2025-26-01.
Acknowledgement
This research was supported by The VNUHCM-University of Information Technology’s Scientific Research Support Fund.

### v2

@@ 6.3 Evaluation on Complex Real-World Exploitation Chains
6.3.1 Performance without RAG (No-RAG)
@@
To assess the baseline capabilities of our proposed system xOffense in realistic offensive security scenarios, we conducted experiments on the same set of six real-world vulnerable machines as utilized in the VulnBot [17] evaluation: Victim1, Library2, Sar, WestWild, Symfonos2, and Funbox. This machine set, originally derived from the AI-Pentest-Benchmark, covers a diverse range of exploitation challenges, including misconfigurations, weak authentication, remote code execution, privilege escalation, and multi-step attack chains. By adopting this identical set, we ensure methodological consistency and enable a direct, fair comparison with prior work.
@@
The experiments were conducted in a fully autonomous mode, without any human intervention or RAG support. Each target machine was tested in five independent runs, and the reported performance represents the best sub-task completion rate per machine, following the AI-Pentest-Benchmark scoring methodology. Figure 3 presents the comparative results across multiple models, including VulnBot-Llama3.1-405B, VulnBot-DeepSeek-v3, their respective base models, and our proposed Qwen3-32B variants (base and finetuned).
@@
The results reveal several noteworthy patterns. First, Qwen3-32B-finetune consistently surpasses its base counterpart across all six machines, with particularly significant improvements on Victim1 (+0.55), Library2 (+0.30), and WestWild (+0.63). These gains highlight the effectiveness of domain-specific fine-tuning in strengthening the model’s exploitation reasoning and procedural robustness. Second, while VulnBot-DeepSeek-v3 remains highly competitive, achieving the highest score on Victim1 (0.83) and WestWild (0.71), our fine-tuned Qwen3-32B achieves comparable or superior performance on most other machines, including leading results on Sar (0.58) and Funbox (0.54).
@@
Notably, performance disparities are strongly correlated with the complexity of exploitation chains. Targets such as Symfonos2 and Funbox, which demand multi-stage privilege escalation and exploitation of non-trivial service configurations, clearly benefit from the enhanced contextual reasoning introduced via fine-tuning. This observation underscores the critical role of model specialization in addressing the inherent unpredictability and diversity of real-world pentest environments. In summary, the No-RAG evaluation confirms that xOffense-Qwen3-32B-finetune can autonomously achieve competitive, and in some cases state-of-the-art, performance in realistic offensive security scenarios, even without external retrieval augmentation. This establishes a robust performance baseline for subsequent RAG-enhanced evaluations.
@@
Figure 3: Comparison of sub-task completion rates across six real-world vulnerable machines in a No-RAG setting.
@@
Figure 4: Comparison of sub-task completion rates across six real-world vulnerable machines with RAG setting.
6.3.2 Performance with RAG (RAG)
@@
When augmenting the evaluation with the Knowledge Repository module, a substantial shift in performance trends emerges across the six real-world exploitation targets (see Fig. 4). Compared to the baseline (No-RAG), the Qwen3-32B-Finetune model demonstrates marked improvement, achieving perfect completion scores on Victim1 and WestWild (1.00) and notable gains on Library2 (+0.20) and Symfonos2 (+0.16). Similarly, moderate increases are observed for Sar and Funbox, reflecting the model’s enhanced capability to navigate multi-step attack chains when supported by targeted, contextually relevant prior knowledge.
@@
The gains are less pronounced for Qwen3-32B-Base, with performance remaining comparatively low on challenging targets such as Symfonos2 (0.13) and WestWild (0.25). This disparity underscores the role of fine-tuning in maximizing the benefits of retrieval augmentation, as without alignment to domain-specific exploitation strategies, the retrieved information alone is insufficient to ensure consistent execution success. When compared against VulnBot baselines, Qwen3-32B-Finetune with RAG achieves competitive or superior results in four out of six targets, matching the best baseline performance on Library2 (0.80) and surpassing it on Victim1 and WestWild. This suggests that RAG integration not only mitigates the limitations of the base model but also allows the fine-tuned variant to close the gap, or in certain scenarios, outperform human-assisted frameworks.
@@
These improvements can be attributed to three key factors: (1) the retriever’s ability to surface high-relevance exploitation procedures from a curated cybersecurity corpus, (2) the fine-tuned model’s capacity to integrate external information into coherent multi-step reasoning, and (3) the reduction of hallucination-driven dead ends, which are particularly detrimental in constrained exploitation environments. Collectively, these findings reinforce the notion that RAG is a critical enabler for scalable, high-fidelity automated pentest in complex real-world settings.
@@ 7.1 Internal validity
The fine-tuning of Qwen3-32B on a CoT-enriched pentest dataset introduces potential internal threats. Certain vulnerability classes and exploitation strategies are disproportionately represented, which may bias the model toward specific attack vectors while limiting its capacity to generalize to underrepresented scenarios. Moreover, the integration of prompting strategies and toolchains may embed implicit task-specific heuristics, raising the possibility that reported improvements partly reflect dataset artifacts rather than genuine reasoning ability. Such factors must be considered when interpreting performance gains on structured benchmarks.
@@ 7.2 External validity
The evaluation settings of AutoPenBench and AI-Pentest-Benchmark, which approximate realistic penetration workflows, yet cannot fully capture the heterogeneity of production-scale environments. Operational networks often exhibit greater variability in topology, non-standard configurations, active defenses, and deception mechanisms that remain absent from current benchmarks. In addition, adversarial tactics evolve over time, whereas benchmarks are necessarily static. Consequently, the generalizability of results to enterprise systems, heterogeneous infrastructures, or zero-day exploitation scenarios should be regarded with caution.
@@ 7.3 Construct validity
Task completion rate and exploitation success were employed as primary evaluation metrics. While suitable for quantifying functional effectiveness, these measures neglect other dimensions that are central to pentest practice. Attributes such as stealth, efficiency of resource utilization, time-to-compromise, and resilience against detection are critical to operational realism yet remain unaccounted for in the adopted benchmarks. Furthermore, binary success measures fail to capture partial progress or incremental compromise, potentially obscuring nuances in agent behavior across complex exploitation chains.
@@ 7.4 Reliability
In sum, although the reported findings provide strong evidence of the capabilities of xOffense, these validity concerns underscore the need for broader empirical validation. Expanding evaluations to encompass more diverse infrastructures, adversarially adaptive defenses, and richer performance metrics would strengthen the robustness, scalability, and practical applicability of autonomous pentest systems.
8 Ethical Considerations and Responsible Use
This work addresses autonomous penetration testing, a domain with inherent dual-use implications. While the proposed xOffense framework is intended to support defensive cybersecurity practices, it may also be subject to misuse if deployed without appropriate safeguards. The system is designed strictly for authorized security assessment, research, and training purposes in controlled environments, such as benchmark platforms and laboratory settings, and must only be used with explicit legal authorization from system owners. It is not intended for unauthorized exploitation of real-world systems.
Due to its capability to automate multi-stage penetration testing workflows, the framework may lower the barrier for conducting offensive operations, including large-scale vulnerability discovery and exploitation. This highlights the importance of responsible usage and user awareness when interacting with such systems. Users are expected to adhere to applicable legal frameworks, institutional policies, and professional ethical standards when deploying or extending the proposed approach.
To mitigate potential risks, all experiments in this study are conducted strictly within sandboxed and publicly available benchmark environments, including AutoPenBench and AI-Pentest-Benchmark, without interaction with live production systems or undisclosed vulnerabilities. In practical deployments, additional safeguards are necessary, including restricting execution to authorized and monitored environments, enforcing logging and auditing of system actions, limiting tool usage and external network access, and incorporating human-in-the-loop validation for high-impact decisions. Future implementations should further integrate policy-based constraints and access control mechanisms to prevent unintended or malicious usage.
Beyond its offensive capabilities, this line of research also motivates the development of more robust defensive mechanisms. Understanding how autonomous agents identify and exploit vulnerabilities can inform the design of improved intrusion detection systems, adaptive defenses, and secure system architectures. As such, this work contributes not only to offensive security automation but also to advancing defensive cybersecurity research.
This research adheres to responsible cybersecurity and AI research practices. The study does not involve zero-day exploitation or unauthorized system access, and all evaluations are conducted on simulated or explicitly permitted targets. Future extensions should follow coordinated disclosure principles and comply with relevant legal and ethical standards, in line with established guidelines for responsible AI and cybersecurity research.
@@ 9 Conclusion and Future Work
This work presented xOffense, an independent, fully autonomous multi-agent framework for pentest, designed to address persistent limitations in existing systems such as context loss, limited reasoning continuity, and dependence on large proprietary models. By integrating a fine-tuned, mid-scale open-source LLM (Qwen3-32B) with a novel grey-box phase prompting mechanism and a purpose-built orchestration architecture, xOffense achieves accurate multi-stage decision-making and robust tool integration across the entire pentest lifecycle.
Our evaluation on AutoPenBench and AI-Pentest-Benchmark demonstrated that xOffense consistently outperforms both larger commercial LLMs, such as GPT-4o and LLaMA3-70B, and leading open-source baselines, such as PentestGPT and VulnBot-LLaMA3-405B. The framework attained an overall task completion rate of 72.72% and a sub-task completion rate of up to 79.17%, while successfully exploiting complex real-world targets. These results highlight that a domain-specialized, mid-scale model, when paired with targeted reasoning guidance, can match or exceed the capabilities of state-of-the-art large-scale systems, offering a cost-effective and reproducible solution for autonomous offensive security operations.
Future work will explore three main directions. First, we aim to optimize the command generation module, potentially through structured function calling, to further improve execution precision. Second, we plan to enhance the robustness of long-running process handling and strengthen the retrieval-augmented generation mechanism with automated updates from vulnerability intelligence sources such as ExploitDB and GitIngest. Third, we intend to extend xOffense’s capabilities to support advanced web and GUI interactions via browser automation, enabling it to tackle a broader range of pentest scenarios.